OpenStudy (anonymous):

7. Solve: |x + 1| = 2 { -3 , 3 } { 3 , -1 } { 3 , 1 } { -3 , 1 }

hartnn (hartnn):

when |a|=b a=b OR a=-b

hartnn (hartnn):

|x+1|=2 will be x+1 =2 OR x+1=-2 solve these, u get 2 values of x

OpenStudy (anonymous):

it jus says solve

hartnn (hartnn):

yeah, solve these 2 : x+1 =2 OR x+1=-2 x+1= 2 whats x from here ?

OpenStudy (anonymous):

i dunno

hartnn (hartnn):

x+1=2 what number when added to 1 , gives 2 ?

OpenStudy (anonymous):

. Multiply the binomials: (2x + 1) (3x - 3) A. 25x2 - 15x - 16 B. 6x2 - 3x - 3 C. 25x2 + 16 D. 25x2 + 20x - 16 help me with this pleasee i need an answer

hartnn (hartnn):

that is simpler. what is 3x * 2x = ?

hartnn (hartnn):

or simply what is 3*2 = ?

OpenStudy (anonymous):

the 2 is an exponent

hartnn (hartnn):

i know, 3*2 = ?